OpenStudy (anonymous):

the equation is y=5x-5 complete the table. x-2 -1 0 1 2 y--------------

OpenStudy (anonymous):

Just substitue your x-values into the equation you have and solve for y. Example for x = 2: y = 5(2) - 5 y = 5 Just do this for the rest :D

OpenStudy (anonymous):

can you help me step by step

OpenStudy (anonymous):

Surely :) Okay, so we have 5 x-values: -2, -1, 0, 1, 2 We simply put each into the equation y = 5x-5 to find y for each.

OpenStudy (anonymous):

So x = -2, we replace x with -2 in the equation: y = 5(-2) - 5 y = -10 - 5 y = -15

OpenStudy (anonymous):

so-2 is 15

OpenStudy (anonymous):

-2 is -15

OpenStudy (anonymous):